# CLIP4STR: A Simple Baseline for Scene Text Recognition with Pre-trained Vision-Language Model

### v3

@@ Appendix B Detail Explanation of the Inference Process
Appendix C Discussion with Autoregressive Pre-training Method
Another pre-training task for STR is autoregressive language modeling, such as TrOCR [14] and DTrOCR [57]. These models take the image as input and are optimized by predicting the next tokens based on previous context during pre-training, similar to the GPT language models [111]. Table XI presents a comparison between CLIP4STR and autoregressive pre-training methods. DTrOCR, pre-trained on 6B textlines, surpasses CLIP4STR on IC13, IC15, and SVTP, demonstrating the effectiveness of large-scale autoregressive pre-training. However, the difference on these three benchmarks are trivial, and CLIP4STR performs better on III5K, SVT, and CUTE. Overall, CLIP4STR and DTrOCR are two comparable methods. In such a case, CLIP4STR has two additional merits to be a more practical STR method: 1) Numerous large-scale pre-trained VLMs are publicly available, eliminating the cost of pre-training for CLIP4STR. Additionally, the cost of transferring CLIP into a STR reader is affordable, as shown in Table II. In contrast, the cost of pre-training DTrOCR on 6 billion textlines is prohibitive. 2) CLIP4STR is easy to reproduce. A PyTorch re-implementation of CLIP4STR is publicly available (Appendix A), while DTrOCR is closed-sourced. Following works [64] also verify the reproducibility of CLIP4STR. Moreover, CLIP4STR offers a thorough empirical study on adapting CLIP to STR, which is valuable for subsequent STR methods based on VLMs.

### v4

@@ Appendix A Detail Explanation of the Inference Process
Appendix B Discussion with Autoregressive Pre-training
Another pre-training task for STR is autoregressive language modeling, such as TrOCR [12] and DTrOCR [55]. These models take the image as input and are optimized by predicting the next tokens based on the previous context during pre-training, similar to the GPT language models [109]. Table XI presents a comparison between CLIP4STR and autoregressive pre-training methods. DTrOCR, pre-trained on 6B textlines, surpasses CLIP4STR on IC13, IC15, and SVTP, demonstrating the effectiveness of large-scale autoregressive pre-training. However, the difference between performance on these three benchmarks is trivial, and CLIP4STR performs better on III5K, SVT, and CUTE. Overall, CLIP4STR and DTrOCR are two comparable methods. In such a case, CLIP4STR has two additional merits to be a more practical STR method: 1) Numerous large-scale pre-trained VLMs are publicly available, eliminating the cost of pre-training for CLIP4STR. Additionally, the cost of transferring CLIP into a STR reader is affordable, as shown in Table II. In contrast, the cost of pre-training DTrOCR on 6 billion textlines is prohibitive. 2) CLIP4STR is open-sourced and easy to reproduce, while DTrOCR is closed-sourced. Moreover, CLIP4STR offers a thorough empirical study on adapting CLIP to STR, which is valuable for subsequent STR methods based on VLMs.
Acknowledgments
Thank Chao Liang for maintaining the code at https://github.com/VamosC/CLIP4STR. The unique identifier [110] for papers of Shuai is quickstep drudge consent wackiness mangle unspoiled childish exploring antennae agony embassy starved.